what is the distributive property to 6(501)
Found 2 solutions by stanbon, jsmallt9:
Answer by stanbon(75887)

what is the distributive property to 6(501)
-----------------------------------
6(6*100 + 0*10 + 1) = 36*100 + 0 +6 = 3606
==============================================
Cheers,
Stan H.

Answer by jsmallt9(3759)

I think the problem is to use the Distributive Property to multiply 6(501). As I hope you'll see, using the Distributive Propoerty to multiply when one of the numbers is near a multiple of power of ten (like 501).

Start by rewriting the number near a multiple of a power of ten as that multiple of a pwer of ten plus or minus some small number:
6(501) = 6(500 + 1)
Next use the Distributive Property to multiply:
6(501) = 6(500 + 1) = 6*500 + 6*1
The two multiplications should be pretty easy. Often they are ones you can do in your head:
6(501) = 6(500 + 1) = 6*500 + 6*1 = 3000 + 6 = 3006

So even though there are two multiplications and an addition this way it is often easier and faster than "normal" multiplication (when one of the numbers is near a multiple of a power of ten).